OpenStudy (anonymous):

calculate the average rate of change for the function g(x)= 4x^2 + x -1 over the interval 0.5≤x≤1

OpenStudy (shubhamsrg):

Any attempt ?

OpenStudy (anonymous):

i substituted the values of 1 and o,5 into the equation and then used the slope equation to get the answer.... but i kept getting it wrong :S

OpenStudy (shubhamsrg):

Can you show your work ?

OpenStudy (anonymous):

4(1)² +1-1 =4(1)+1-1 =4+1-1 =4 4(0.5)²+0.5-1 =4(.25)+0.5-1 =1+0.5-1 =0.5 (4-0.5)/(1-0.5) .... OH ! LOL I KNOW WHERE I WENT WRONG WITH MY EQUATION NOW!

OpenStudy (anonymous):

WHOOPS...... sorry ._.